Let a be the position vector of a given point (x1, y1, z1), and r the position vector of any point (x, y, z). Describe the locus of r if (a) I r - a I = 3, (b) (r-a). a = 0, (c) (r-a).r = 0.

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers
Aditya Gupta
4 years ago
(a) | r - a | = 3 is a sphere with centre (x1, y1, z1) and radius 3.
(b) (r-a). a = 0 is a plane passing through (x1, y1, z1) with normal a.
(c) (r-a).r = 0 is a sphere with centre (x1/2, y1/2, z1/2) and radius |a|/2.